OpenStudy (anonymous):

Add or subtract the monomials. 6 - 12p^5 - 3p + 8 - 8p^5

jimthompson5910 (jim_thompson5910):

Combine like terms So for example: \[\Large -12p^5 - 8p^5 = (-12-8)p^5 = -20p^5\] So \[\Large -12p^5 - 8p^5=-20p^5\]